I have a chance to buy a 23 inch or 24 inch diagonal monitor for my computer system. Each monitor has a 16:10 ratio (Wide screen). What are the measurements (L, W, and Area) of each monitor? Which has the biggest area?
The only thing I could think of to try would be to use the pythagorean theorem as a first step, but then that leaves out the ratio part of the question, so I think that may be wrong. We have recently studied quadratic equations, so maybe I'm supposed to put it in that form to solve? I'm really lost!

Answer by stanbon(75887)

I have a chance to buy a 23 inch or 24 inch diagonal monitor for my computer system. Each monitor has a 16:10 ratio (Wide screen).
What are the measurements (L, W, and Area) of each monitor?
--------
24-in diagonal:
(16x)^2 + (10x)^2 = 24^2
356x^2 = 576
x = 1.272
----
width= 16x = 20.35 inches
length = 10x = 12.72 inches
Area = 258.88 sq. in.
------------------------------------------
Do the same for a diagonal of 23 inches.
----
Which has the biggest area?
=======================================
Cheers,
Stan H.